Changes between Version 1 and Version 2 of Ticket #615, comment 2


Ignore:
Timestamp:
Jun 6, 2017 6:10:31 PM (23 months ago)
Author:
lah
Comment:

Legend:

Unmodified
Added
Removed
Modified
  • Ticket #615, comment 2

    v1 v2  
    99* There were some files without dark frames in the morning, but these were repeated, so not a problem. They are in the reject_files directory. These shouldn't be a problem in the afternoon, but if something goes wrong check for this.
    1010* Although the data were acquired in one day the 2 sets are named d-1 and d-2 for consistency. They may be renamed.
    11 * Since adding O back in 2 lines have failed so will need checking. Make sure all the lamps are correct in the spectral lines file you use and rerun with different offsets. The scripts now automatically ignore ridiculous FWHM but there may still be a ~10 nm wide peak in the VNIR that shouldn't be there.
     11* Since adding O back in 2 lines have failed so will need checking (O lamp may not be reliable as the signal was faint - check ). Make sure all the lamps are correct in the spectral lines file you use and rerun with different offsets. The scripts now automatically ignore ridiculous FWHM but there may still be a ~10 nm wide peak in the VNIR that shouldn't be there.
    1212* Should check with Chris that the method implemented for Mylar is correct before using.
    1313* For the data quality report you can compare results with Specim's instead of last year's. Their reports are in /users/rsg/arsf/calibration/loan_fenix_2017